Parth (parthkohli):

Is this a valid mathematical sentence?\[\rm f:3\mapsto 9 ~where~f(x) = x^2 \]

OpenStudy (amistre64):

i can read it, so its prolly not ;)

Parth (parthkohli):

Is this a better way to write it?\[\rm f:x\mapsto x^2 \implies f:3\mapsto9 \]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

That's invalid notation.\[(f=x\mapsto x^2)\implies(f(3)=9)\]

OpenStudy (anonymous):

f:{3} to {9}, x maps to x^2 is a function definition (to which u may sometimes add the codomain as well).

OpenStudy (anonymous):

any function is defined over sets if you enclose the 3 and 9 in parenthesis and say "this set is my domain and the other one is my co-domain", then its fine

OpenStudy (anonymous):

it can also be f(x) = 3x we can write in many ways we can not derive a function from a single relation

Parth (parthkohli):

What I want to write is: 3 maps to 9 in the function f(x) = x^2

Parth (parthkohli):

What'd be a rather good way to do so?

OpenStudy (anonymous):

write f(3)=9

Parth (parthkohli):

No, I'm trying to use much of the mapping symbol to get the hang of it.

OpenStudy (anonymous):

f:{3}->{9} such that f(x)=x^2

OpenStudy (anonymous):

Use Arrow Diagram

OpenStudy (anonymous):

f: Set to ( arrow) Set, x maps to (arrow with tail) f(x) defines a function.

OpenStudy (unklerhaukus):

?

Parth (parthkohli):

So\[\rm f:3\mapsto9|f(x) = x^2\]

OpenStudy (anonymous):